# FLORES: A Reconfigured Wheel-Legged Robot for Enhanced Steering and Adaptability

### v1

@@ Abstract
Wheel-legged robots integrate the agility of legs for navigating rough terrains while harnessing the efficiency of wheels for smooth surfaces. However, most existing designs do not fully capitalize on the benefits of both legged and wheeled structures, which limits overall system flexibility and efficiency. We present FLORES (reconfigured wheel-legged robot for enhanced steering and adaptability), a novel wheel-legged robot design featuring a distinctive front-leg configuration that sets it beyond standard design approaches. Specifically, FLORES replaces the conventional hip-roll degree of freedom (DoF) of the front leg with hip-yaw DoFs, and this allows for efficient movement on flat surfaces while ensuring adaptability when navigating complex terrains. This innovative design facilitates seamless transitions between different locomotion modes (i.e., legged locomotion and wheeled locomotion) and optimizes the performance across varied environments. To fully exploit FLORES’s mechanical capabilities, we develop a tailored reinforcement learning (RL) controller that adapts the Hybrid Internal Model (HIM) with a customized reward structure optimized for our unique mechanical configuration. This framework enables the generation of adaptive, multi-modal locomotion strategies that facilitate smooth transitions between wheeled and legged movements. Furthermore, our distinctive joint design enables the robot to exhibit novel and highly efficient locomotion gaits that capitalize on the synergistic advantages of both locomotion modes. Through comprehensive experiments, we demonstrate FLORES’s enhanced steering capabilities, improved navigation efficiency, and versatile locomotion across various terrains. The open-source project can be found at https://github.com/ZhichengSong6/FLORES-A-Reconfigured-Wheel-Legged-Robot-for-Enhanced-Steering-and-Adaptability.git.
@@ I INTRODUCTION
Wheel-legged robots have emerged as a promising solution for diverse locomotion tasks, which seamlessly integrate the agile walking capabilities of legged robots for navigating rough terrain with the high-efficiency motion of wheeled robots on flat surfaces [1, 2, 3, 4]. These hybrid systems have shown remarkable potential in applications such as exploration [5], rescue [6], and delivery [7]. However, despite these significant advancements, two fundamental challenges persist in fully capitalizing on the advantages of both legged and wheeled locomotion: wheel-leg mechanical design and control strategies.
@@
The structural design of current wheel-legged robots generally falls into two distinct categories, each with its own set of limitations. The first category treats the legged structure as an active suspension system. This design often results in fewer actuators [8] or limited joint ranges [9], restricting the robot’s performance in environments that require rapid or dynamic mode switching. Some designs within this category, such as [10], utilize actuators that provide low output speeds, while others incorporate harmonic reducers [11], which offer high torque at the expense of speed.
Combined with a heavy leg structure, this makes the system unsuitable for high-dynamic hybrid locomotion. The second category involves directly modifying existing legged platforms by incorporating wheels [3, 4]. In this configuration, the wheel serves as a supplementary feature rather than a core design element. While it enhances straight-line locomotion, it proves less adaptable for more intricate movements, such as turning or traversing cluttered environments [12, 13, 14, 15], as most complex maneuvers beyond straight-line motion require leg lifting and repositioning, during which the increased weight of the wheels consistently imposes a negative impact on the robot’s agility and energy efficiency, particularly when navigating unstructured or rough terrains.
@@
We propose FLORES, an innovative wheel-legged robot that differs from conventional design approaches. As shown in Fig. 2, we reconfigure the front leg structure by incorporating a hip-yaw joint, which significantly enhances the robot’s steering capabilities. This new configuration integrates the wheel-leg assembly into a cohesive system, combining the hip-yaw, hip-pitch, knee-pitch, and wheel joints. The hip-yaw joint acts as a comprehensive steering module, enabling efficient locomotion on relatively flat surfaces without the need for frequent leg lifting, similar to the steering mechanism of a vehicle. This design improves both steering capabilities and adaptability in complex environments while maintaining the efficiency of linear motion. Moreover, the unmodified rear legs preserve the robot’s ability to navigate and stabilize on more complex, unstructured terrains, ensuring versatility across various environments. To optimize the capabilities of this hybrid configuration, we modify the Hybrid Internal Model (HIM) framework [23] to align with our distinct mechanical design. We customize the reward structure tailoring to FLORES’s innovative features, which places particular emphasis on steering, dynamic adaptability, and the ability to navigate complex terrains. As demonstrated in our experiment, our method enables dynamic optimization of locomotion across a wide range of challenging scenarios, showcasing the synergy between FLORES’s innovative design and its control framework.
@@
  • Demonstration of Superior Performance: We demonstrate FLORES’s superior steering abilities, increased navigation efficiency, and adaptability for locomotion across a range of terrains.
@@ II-A Wheel-Legged Robot Platforms
Wheels with Active Suspension. Bipedal designs such as Ascento [9] provide efficient movement on flat terrain, but due to their limited number of actuators and restricted joint range, they can hardly walk and are primarily designed for jumping to overcome obstacles. On the other hand, quadrupedal designs like Momaro [24] and Centauro [11, 25, 26] can simultaneously drive and walk. Nevertheless, Centauro’s performance is constrained by harmonic reducers in its actuators and heavy leg structures, which limit its capacity for high-dynamic locomotion. While harmonic reducers provide high torque, they suffer from low output speeds and poor resistance to continuous impacts, reducing the robot’s agility and overall performance in dynamic environments. Although series elastic actuation is employed to enhance structural integrity, it introduces significant nonlinearity in motor control. Similarly, Momaro utilizes servo motors as actuators, which also deliver low output speeds.
Wheels as Add-on Features. The DRC-HUBO+ [27] is a bipedal robot that incorporates an additional wheel structure. However, it struggles with hybrid locomotion, as reconfiguration is necessary when switching between walking and driving modes. A more advanced design is ANYmal [3, 4], which operates in walking, driving, or hybrid modes. Built on the legged version of ANYmal, a state-of-the-art quadrupedal robot platform, this design offers improved versatility. Nonetheless, the wheel structure primarily benefits forward straight motion, thereby fully utilizing the advantages of wheeled locomotion. In complex environments, however, other essential movements such as turning and lateral movement, which require frequent foot lifting, result in reduced energy efficiency.
@@ II-B Control Strategies for Wheel-Legged Robot
Learning-based Control. Recent advances in massively parallel RL simulations [30] have significantly reduced the time required for policy development, making RL increasingly appealing for tasks that involve diverse terrains or unpredictable obstacles, where traditional model-based approaches may struggle to maintain optimal performance. For instance, ANYmal employs a recurrent neural network-based policy to facilitate robust and versatile locomotion [12]. Additionally, Whleaper, a bipedal wheel-legged robot, applies the Proximal Policy Optimization (PPO) algorithm to facilitate reliable walking and jumping locomotion [21]. Similarly, PPO with an asymmetric actor-critic architecture enables blind stair climbing for various wheel-legged robots [22]. In our work, we adopt the HIM framework [23], which integrates classical Internal Model Control (IMC) principles with the PPO algorithm. This framework estimates environmental disturbances, such as ground friction and terrain elevation, using only the robot’s proprioceptive data from inertial measurement units (IMUs) and joint encoders, thereby eliminating the need for additional sensors.
@@ III-A Hardware Design
FLORES is designed to achieve versatile locomotion in various terrains, fully exploiting the advantage of both legged and wheeled structures [^1] All the hardware design and a datasheet for all electronics can be found in our project website..
@@
Mechanical Design. FLORES has a total of 16 degrees of freedom (DoFs) with 4 DoFs per leg, as shown in Fig. 2(a). Each leg is equipped with 3 motors near the torso: two motors provide hip DoFs (hip-yaw & pitch for front legs/ hip-roll & pitch for rear legs), and one for actuating the knee-pitch DoF. Additionally, a motor is mounted at the end of each leg to drive the wheel. The hip-pitch and knee-pitch motors are co-axially located and the knee-pitch DoF is actuated through a parallel four-bar linkage mechanism, which helps reduce leg inertia, and all other DoFs are directly actuated by motors.
In particular, as demonstrated in Fig. 2(b), the front legs feature a hip-yaw DoF, which acts as a steering module and significantly improves the robot’s maneuverability on unstructured terrain. To obtain enough steering space for front legs, the distance between front hip joints is larger than the back hip joints, which can prevent the leg collide with the body when a big heading angle is needed. The hip-yaw DoF has a range of motion from -35° to 100°, providing greater flexibility in adjusting the robot’s heading. This wide range enhances agility, allowing the robot to quickly change directions—especially beneficial in tight or complex environments. Meanwhile, as shown in Fig. 2(c), the rear legs with their typical hip-roll configuration maintain the necessary stability to navigate unstructured terrains, ensuring the robot’s adaptability in various environments. As such, each front leg is equipped with motors for hip-yaw, hip-pitch, knee-pitch, and wheel actuation, while the rear legs retain the typical hip-roll configuration.
@@
Actuators. Brushless DC motors with planetary reducers from Motorevo are used to actuate all the joints. Specifically, all motors feature a 10:1 reduction ratio. The hip-yaw, hip-roll, hip-pitch, and knee-pitch motors have a maximum continuous torque of 32 N·m, while the wheel motors have a maximum continuous torque of 8 N·m. Each motor is equipped with a single encoder to provide position, velocity, and current measurement feedback. Motor control is achieved via the EtherCAT communication protocol. The relatively low reduction ratios of planetary reducers are chosen for two primary reasons: first, they facilitate high-speed motion, which is essential for dynamic locomotion; second, these quasi-direct drive motors provide relatively accurate torque feedback through current measurements, eliminating the need for additional sensors. Furthermore, planetary reducers are capable of withstanding significant unexpected impacts, which frequently occur during dynamic movement.
Manufacturing. The body plates and legs of the robot are constructed from off-the-shelf aluminum 6061, while certain electronic mounts are 3D-printed using Polylactic Acid (PLA). Additionally, a grid of screw holes is incorporated between the front motors, allowing for the installation of sensors such as lidar and cameras to facilitate future developments.
@@ III-B1 Overview
As illustrated in Fig. 3, the software architecture is composed of two sequential phases to realize robust locomotion control on our wheel-legged robot. In Phase $\mathrm{I}$, training is conducted in simulation using Isaac Gym [30], where the robot’s state is described by joint velocity, joint position, angular velocity, gravitational force, and the previous action. These inputs are processed by the HIM, which integrates an encoder-based state representation with a customized reward function. The policy network is then optimized via PPO, providing target wheel velocities and joint positions as output commands. In Phase $\mathrm{II}$, the learned policy is first validated in MuJoCo [31], allowing for refined fine-tuning under more realistic dynamics. Subsequently, the final policy is deployed onto the real robot. Hardware feedback—including signals from the IMU and motor encoders are continuously fed into the control policy, which runs on an embedded compute unit. By continuously sending real-time torque commands, the robot achieves smooth sim-to-real transitions and robust performance in physical environments.
@@ III-B2 Reinforcement Learning for FLORES
These combined inputs constitute the observation $\mathbf{S}_{t}$, which is then processed by the control policy to generate the appropriate control signals.
Actions. The action $\mathbf{a}_{t}\in\mathbb{R}^{16}$ contains two different categories of actions: $\mathbf{a}^{\text{leg}}_{t}\in\mathbb{R}^{12}$ and $\mathbf{a}^{\text{wheel}}_{t}\in\mathbb{R}^{4}$. $\mathbf{a}^{\text{leg}}_{t}$ is the bias between the desired joint angle and the default joint angle for 12 joints, except for the wheels. A fixed action scale ${k}$ is multiplied by the action to reduce the instability of the network output. The desired joint angle is the sum of the default joint angle and the product of the scale factor and the action, which can be defined as
@@
Reward. Table I presents the reward functions for our robot. In this context, $e_{v}$ and $e_{\omega_{z}}$ represent the tracking errors for linear velocity and angular velocity, respectively, while the function $n(e)$ indicates whether the error $e$ is close to zero. The variables $v_{z}$, $\omega_{x}$, $\omega_{y}$, and $g^{b}_{xy}$ denote the base velocity along the $z$-axis, angular velocity around the $x$-axis, angular velocity around the y-axis, and the projected gravity in the $x$ and $y$ directions, respectively. Additionally, $h$, $h_{d}$ and $\mathcal{I}$ represent the height of the robot, the target height of the robot, and an indicator function that equals zero when the robot is in motion. The variable $\tau$ refers to the output torque of all the joints, while $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$, $\sigma_{4}$, and $\sigma_{5}$ are user-defined coefficients for the various reward components.
@@
$z$-axis Lin. Vel. $v_{z}^{2}$0.1
$x$-axis & $y$-axis Ang. Vel. $\omega_{x}^{2}+\omega_{y}^{2}$0.05
Orientation $(\mathbf{g}_{xy}^{b})^{2}$0.2
@@
Joint Acc. $\left\|\ddot{\mathbf{q}}\right\|_{2}^{2}$2.5e-7
Joint Power $|\bm{\tau}||\dot{\mathbf{q}}|^{T}$5e-5
Torques $|\bm{\tau}|^{2}$5e-5
Action Rate $(\mathbf{a}_{t-1}-\mathbf{a}_{t})^{2}$0.01
Smoothness $(\mathbf{a}_{t}-2\mathbf{a}_{t-1}+\mathbf{a}_{t-2})^{2}$0.01

### v2

# A Reconfigured Wheel-Legged Robot for Enhanced Steering and Adaptability

@@ Abstract
Wheel-legged robots integrate leg agility on rough terrain with wheel efficiency on flat ground. However, most existing designs do not fully capitalize on the benefits of both legged and wheeled structures, which limits overall system flexibility and efficiency. We present FLORES, a novel wheel-legged robot design featuring a distinctive front-leg configuration that sets it beyond standard design approaches. Specifically, FLORES replaces the conventional hip-roll degree of freedom (DoF) of the front leg with hip-yaw DoFs, and this allows for efficient movement on flat surfaces while ensuring adaptability when navigating complex terrains. This innovative design facilitates seamless transitions between different locomotion modes (i.e., legged locomotion and wheeled locomotion) and optimizes the performance across varied environments. To fully exploit FLORES’s mechanical capabilities, we develop a tailored reinforcement learning (RL) controller that adapts the Hybrid Internal Model (HIM) with a customized reward structure optimized for our unique mechanical configuration. This framework enables the generation of adaptive, multi-modal locomotion strategies that facilitate smooth transitions between wheeled and legged movements. Furthermore, our distinctive joint design enables the robot to exhibit novel and highly efficient locomotion gaits that capitalize on the synergistic advantages of both locomotion modes. Through comprehensive experiments, we demonstrate FLORES’s enhanced steering capabilities, improved navigation efficiency, and versatile locomotion across various terrains. The open-source project can be found at https://github.com/ZhichengSong6/FLORES.
@@ I INTRODUCTION
Wheel-legged robots have emerged as a promising solution for diverse locomotion tasks, which seamlessly integrate the agile walking capabilities of legged robots for navigating rough terrain with the high-efficiency motion of wheeled robots on flat surfaces [10, 13, 2, 4]. These hybrid systems have shown remarkable potential in applications such as exploration, rescue, and delivery [29, 25]. However, despite these significant advancements, two fundamental challenges persist in fully capitalizing on the advantages of both legged and wheeled locomotion: wheel-leg mechanical design and control strategies.
@@
The structural design of current wheel-legged robots generally falls into two categories, each with its own limitations. The first category treats the legged structure as an active suspension system. This design often results in fewer actuators [24] or limited joint ranges [16], restricting the robot’s performance in environments that require rapid or dynamic mode switching. Representative systems such as SherpaTT further demonstrate actively articulated suspension designs in Mars-analog field tests [7]. Some designs within this category, such as [27], utilize actuators that provide low output speeds, while others incorporate harmonic reducers [11], which offer high torque at the expense of speed. Combined with a heavy leg structure, this makes the system unsuitable for high-dynamic hybrid locomotion.
The second category involves directly modifying existing legged platforms by incorporating wheels [2, 4]. In this configuration, the wheel serves as a supplementary feature rather than a core design element. While it enhances straight-line locomotion, it is less adaptable for more intricate movements, such as turning or traversing cluttered environments [19, 8, 32], as most complex maneuvers beyond straight-line motion require leg lifting and repositioning, during which the increased wheel weight consistently imposes a negative impact on the robot’s agility and energy efficiency, particularly when navigating unstructured or rough terrains.
@@
We propose FLORES, an innovative wheel-legged robot that differs from conventional designs. As shown in Fig. 2, we reconfigure the front leg by incorporating a hip-yaw joint, which significantly enhances the robot’s steering capabilities. The hip-yaw joint acts as a comprehensive steering module, enabling efficient locomotion on relatively flat surfaces without frequent leg lifting, similar to a vehicle steering mechanism. This design improves both steering capabilities and adaptability in complex environments while maintaining the linear-motion efficiency. Moreover, the unmodified rear legs preserve the robot’s ability to navigate and stabilize on more complex, unstructured terrains, ensuring versatility across various environments. To optimize the capabilities of this hybrid configuration, we modify the Hybrid Internal Model (HIM) framework [21] to align with our distinct mechanical design. We customize the reward structure tailoring to FLORES’s innovative features, which places particular emphasis on steering, dynamic adaptability, and the ability to navigate complex terrains. As demonstrated in our experiment, our method enables dynamic optimization of locomotion across a wide range of challenging scenarios, showcasing the synergy between FLORES’s innovative design and its control framework.
@@
  • Demonstration of Superior Performance: We demonstrate FLORES’s superior steering abilities, increased navigation efficiency, and adaptability across diverse terrains, particularly in disturbance-rich and cluttered scenarios requiring frequent reorientation.
@@ II-A Wheel-Legged Robot Platforms
Wheels with Active Suspension. Bipedal designs such as Ascento [16] provide efficient movement on flat terrain, but due to their limited number of actuators and restricted joint range, they can hardly walk and are primarily designed for jumping to overcome obstacles. On the other hand, quadrupedal designs like Momaro [14] and Centauro [11] can simultaneously drive and walk. Nevertheless, Centauro’s performance is constrained by harmonic reducers in its actuators and heavy leg structures, which limit its capacity for high-dynamic locomotion. While harmonic reducers provide high torque, they suffer from low output speeds and poor resistance to continuous impacts, reducing the robot’s agility and overall performance in dynamic environments. Although series elastic actuation is employed to enhance structural integrity, it introduces significant nonlinearity in motor control. Similarly, Momaro utilizes servo motors as actuators, which also deliver low output speeds.
Wheels as Add-on Features. The DRC-HUBO+ [20] is a bipedal robot that incorporates an additional wheel structure. However, it struggles with hybrid locomotion, as reconfiguration is necessary when switching between walking and driving modes. A more advanced design is ANYmal [2, 4], which operates in walking, driving, or hybrid modes. Built on the legged version of ANYmal, a state-of-the-art quadrupedal robot platform, this design offers improved versatility. The wheel structure excels at straight-ahead travel by exploiting pure rolling, which is energy-efficient. However, in complex environments, essential movements like turning and lateral movement that require frequent foot lifting could result in reduced energy efficiency.
@@ II-B Control Strategies for Wheel-Legged Robot
Learning-based Control. Recent advances in massively parallel RL simulations [22] have significantly reduced the time required for policy development, making RL increasingly appealing for tasks that involve diverse terrains or unpredictable obstacles, where traditional model-based approaches may struggle to maintain optimal performance. For instance, ANYmal employs a recurrent neural network-based policy to facilitate robust and versatile locomotion [19]. Additionally, Whleaper, a bipedal wheel-legged robot, applies the Proximal Policy Optimization (PPO) algorithm to facilitate reliable walking and jumping locomotion [33]. Similarly, PPO with an asymmetric actor-critic architecture enables blind stair climbing for various wheel-legged robots [6]. Recent works also explore learning-enhanced disturbance compensation in structured controllers. For example, NeuroMHE [31] combines recurrent networks with moving horizon estimation to infer latent disturbances, while L1-MPC [26] integrates L1 adaptive control into MPC formulations for improved robustness against unmodeled dynamics.
In our work, we adopt the HIM framework [21], which integrates classical Internal Model Control (IMC) principles with the PPO algorithm. This framework estimates environmental disturbances, such as ground friction and terrain elevation, using only the robot’s proprioceptive data from inertial measurement units (IMUs) and joint encoders, thereby eliminating the need for additional sensors.
@@ III-A Hardware Design
FLORES is designed for versatile locomotion across terrains, exploiting both legged and wheeled structures [^1] All the hardware design and a datasheet for all electronics can be found in our project website..
@@
Mechanical Design. FLORES has a total of 16 degrees of freedom (DoFs) with 4 DoFs per leg, as shown in Fig. 2(a). Each leg is equipped with 3 motors near the torso: two motors provide hip DoFs (hip-yaw & pitch for front legs/ hip-roll & pitch for rear legs), and one for actuating the knee-pitch DoF. Additionally, a distal motor drives each wheel. The hip-pitch and knee-pitch motors are co-axially located and the knee-pitch DoF is actuated through a parallel four-bar linkage mechanism, which reduces leg inertia, and all other DoFs are directly actuated by motors. Such parallel linkage mechanisms are commonly adopted in series-parallel hybrid robotic systems for improved structural and dynamic performance [17].
In particular, as demonstrated in Fig. 2(b), the front legs feature a hip-yaw DoF, which acts as a steering module and significantly improves the robot’s maneuverability on unstructured terrain. To obtain enough steering space for front legs, the distance between front hip joints is larger than the back hip joints, which can prevent the leg collide with the body when a big heading angle is needed. The hip-yaw DoF has a range of motion from -35° to 100°, providing greater flexibility in adjusting the robot’s heading. This range improves agility, allowing the robot to quickly change directions—especially beneficial in tight or complex environments. Meanwhile, as shown in Fig. 2(c), the rear legs with their typical hip-roll configuration maintain stability on unstructured terrain, ensuring the robot’s adaptability in various environments. As such, each front leg is equipped with motors for hip-yaw, hip-pitch, knee-pitch, and wheel actuation, while the rear legs retain the typical hip-roll configuration.
@@
Actuators. All joints use Motorevo brushless DC motors with 10:1 planetary reducers. The hip–yaw/roll/pitch and knee–pitch motors provide 32 N·m continuous torque, and the wheel motors 8 N·m. Each motor is equipped with a single encoder to provide position, velocity, and current measurement feedback. Motor control is achieved via the EtherCAT communication protocol. The relatively low reduction ratios of planetary reducers are chosen for two primary reasons: first, they facilitate high-speed motion, which is essential for dynamic locomotion; second, these quasi-direct drive motors provide relatively accurate torque feedback through current measurements, eliminating the need for additional sensors. Furthermore, planetary reducers are capable of withstanding significant unexpected impacts, which frequently occur during dynamic movement.
Manufacturing. The body plates and legs of the robot are constructed from off-the-shelf aluminum 6061, while certain electronic mounts are 3D-printed using Polylactic Acid (PLA).
@@ III-B1 Overview
As illustrated in Fig. 3, the software architecture is composed of two sequential phases to realize robust locomotion control on our wheel-legged robot. In Phase $\mathrm{I}$, training is conducted in Isaac Gym [22], where the robot’s state is described by joint velocity, joint position, angular velocity, gravitational force, and the previous action. These inputs are processed by HIM, which integrates an encoder-based state representation with a customized reward function. The policy network is then optimized via PPO, providing target wheel velocities and joint positions as output commands. In Phase $\mathrm{II}$, the learned policy is first fine-tuned in MuJoCo [30] under more realistic dynamics. Subsequently, the final policy is deployed onto the real robot. Hardware feedback—including signals from the IMU and motor encoders are fed to the policy, which runs on an embedded compute unit. By sending real-time torque commands, the robot achieves smooth sim-to-real transitions and robust performance in physical environments.
@@ III-B2 Reinforcement Learning for FLORES
These inputs constitute the observation $\mathbf{S}_{t}$, which is then processed by the control policy to generate the appropriate control signals.
Actions. The action $\mathbf{a}_{t}\in\mathbb{R}^{16}$ contains two different categories of actions: $\mathbf{a}^{\text{leg}}_{t}\in\mathbb{R}^{12}$ and $\mathbf{a}^{\text{wheel}}_{t}\in\mathbb{R}^{4}$. $\mathbf{a}^{\text{leg}}_{t}$ is the bias between the desired joint angle and the default joint angle for 12 joints, except for the wheels. A fixed action scale k{k} is multiplied by the action to reduce output instability. The desired joint angle is the sum of the default joint angle and the product of the scale factor and the action, which can be defined as
@@
Reward. Our reward design follows a standard Legged-Gym–style shaping [28], with adjustments made solely to the weights for our specific platform. Table I presents the reward functions for our robot. In this context, $e_{v}$ and $e_{\omega_{z}}$ represent the tracking errors for linear velocity and angular velocity, respectively, while the function $n(e)$ indicates whether the error $e$ is close to zero. The variables $v_{z}$, $\omega_{x}$, $\omega_{y}$, and $g^{b}_{xy}$ denote the base velocity along the $z$-axis, angular velocity around the $x$-axis, angular velocity around the y-axis, and the projected gravity in the $x$ and $y$ directions, respectively. Additionally, $h$, $h_{d}$ and $\mathcal{I}$ represent the height of the robot, the target height of the robot, and an indicator function that equals zero when the robot is in motion. The variable $\tau$ refers to the output torque of all the joints, while $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$, $\sigma_{4}$, and $\sigma_{5}$ are user-defined coefficients for the various reward components.
@@
$z$-axis Lin. Vel. $v_{z}^{2}$ -0.1
$x$-axis & $y$-axis Ang. Vel. $\omega_{x}^{2}+\omega_{y}^{2}$ -0.05
Orientation $(\mathbf{g}_{xy}^{b})^{2}$ -1.0
@@
Joint Acc. $\left\|\ddot{\mathbf{q}}\right\|_{2}^{2}$ -2.5e-7
Joint Power $|\bm{\tau}||\dot{\mathbf{q}}|^{T}$ -5e-5
Torques $|\bm{\tau}|^{2}$ -5e-5
Action Rate $(\mathbf{a}_{t-1}-\mathbf{a}_{t})^{2}$ -0.01
Smoothness $(\mathbf{a}_{t}-2\mathbf{a}_{t-1}+\mathbf{a}_{t-2})^{2}$ -0.01